Find the values of p, q and r such that the force

F~ = p y2z3 ˆi − q xyz3jˆ + r xy2z2 ˆk is a conservative force. What is the value of the potential?


1
Expert's answer
2021-02-08T18:37:50-0500

We know that force is conservative if and only if "\vec\nabla \times F\equiv0" (strictly speaking, we also need to suppose that domain of F is simply connected, but in this case it is defined on the whole "\mathbb{R}^3"). Let's calculate "\text{curl }\vec F" :

"\text{curl } \vec F = \begin{pmatrix} \frac{\partial F_z}{\partial y} - \frac{\partial F_y}{\partial z} \\ \frac{\partial F_x}{\partial z} - \frac{\partial F_z}{\partial x} \\ \frac{\partial F_y}{\partial x} - \frac{\partial F_x}{\partial y} \end{pmatrix} \equiv 0"

This condition gives us 3 equations :

  1. "2rxyz^2 + 3 q xyz^2 = 0" that gives "2r+3q=0"
  2. "3py^2 z^2-ry^2 z^2=0" that gives "3p-r=0"
  3. "-qyz^3-2pyz^3=0" that gives "2p+q=0"

This system of equation gives us a family of solutions of a form "q = -2p, r = 3p, p\in \mathbb{R}".

Now to find the potential, we can either "guess" it from the expressions "-\frac{\partial V}{\partial x} = py^2 z^3, -\frac{\partial V}{\partial y} = 2pxyz^3", "-\frac{\partial V}{\partial z} = 3pxy^2 z^2" and thus we get "V = -pxy^2 z^3+const", or for a stricter approach define "V(x,y,z) = V(0) - \int_\gamma \vec{F}(x,y,z) \cdot \vec{\tau} dt", where "\gamma" is any curve joining the origin and the point (x,y,z), "\tau" is the tangent vector to "\gamma" and t is a parametrization of "\gamma". As "\text{curl } \vec F \equiv 0" the choice of "\gamma" doesn't change the integral and thus the definition is consistent. We can choose, for example, a straight line joining the origin with the point (x,y,z) and thus we get ("\gamma(t)=(xt,yt,zt), t\in[0;1], \tau = (x,y,z)" :

"V(x,y,z) = V(0) - p \int_0^1 (xy^2 z^3 t^5+ 2xy^2z^3 t^5+ 3xy^2z^3 t^5)dt=V(0) - pxy^2z^3", we get the same expression as earlier.